# Lumos-1: On Autoregressive Video Generation from a Unified Model Perspective

### v1

@@ 3.1 Spatiotemporal Correlation Injection via MM-RoPE
12:     $\mathcal{C}\leftarrow G_{\phi}(\widetilde{\bm{X}}_{v}^{(t)},AttnMask^{(t)},\mathcal{C})$triangleright Store cache for the $t$-th frame
@@
Moreover, for a model that jointly processes text and visual tokens, the interplay between two modalities is significant to ensure the vision-language alignment. However, the range of positions for representing texts or visual data tends to differ. Contemporary visual generation systems are typically trained with extremely long and descriptive captions [4, 66], despite the low latent resolution of visual data (e.g., a video of resolution $448\times 256\times 25$ becomes $56\times 32\times 7$ after a $8\times 8\times 4$ compression). To balance the two modalities, we propose scaling the 3D positions to ensure a balanced learning. Specifically, we empirically scale the latent 3D positions to the RGB space by multiplying the compression ratio, as shown in figure 5(a). This simple scaling operation, from another perspective, improves the RoPE resolution for visual tokens by slightly accelerating the rotation speed. Experiments in the experiment section demonstrate its efficacy, thereby showing the importance of balancing the two modalities from the RoPE perspective. We acknowledge, however, that given the autoregressive generation nature of videos, this scaling may not be the optimal solution. More advanced and sophisticated solutions are left for future work.
@@ 3.2 Autoregressive Discrete Diffusion Forcing
The most naive generation paradigm (i.e., next-token prediction) suffers from low generation efficiency, rendering it impractical for autoregressive visual generation. In this work, we resort to discrete diffusion [81, 63] to generate visual content, together with temporally causal dependency to enable a temporally autoregressive generation paradigm. However, due to the autoregressive nature of Lumos-1, naive random masking (i.e., a globally random mask) or temporally independent masking (i.e., diffusion forcing [9, 60]) both lead to significant loss imbalance, i.e., visual tokens in the later frames tend to have much lower loss. Since the task difficulty of predicting frames with ample history frame context is considerably easier than predicting the first image given a text prompt or predicting the second frame given the first frame, the model would lead towards optimizing simpler tasks, leading to a degradation of temporal learning.
Training scheme. To resolve this problem, we build upon the basic characteristics of videos–spatial information redundancy. The core problem of imbalanced loss during training stems from spatial information leakage. It is worth noting that the original video diffusion transformer [60] that adopts diffusion forcing does not encounter this issue due to its usage of bidirectional dependency. Facing this challenge, we introduce autoregressive discrete diffusion forcing (AR-DF), a technique that adopts temporal tube masking during the training of autoregressive video generation. Concretely, for every video, we randomly generate a mask pattern for the first frame and then apply this mask pattern repeatedly to later frames in this video. If we denote the multi-modal token sequence $\bm{X}$ composed of text tokens $\bm{X}_{p}$ and visual tokens $\bm{X}_{v}$ and sample a mask ratio $\rho$, the training masking strategy in AR-DF can be formulated as:
@@
$\widetilde{\bm{X}}_{v}^{(t)}\;=\;\bm{M}\odot\bm{X}_{v}^{(t)}\;+\;(1-\bm{M})\,\odot\,[\text{MASK}],\quad t=1,\dots,T$ (8)
@@
[…]
Table 1: Performance comparison on GenEval. In the “#Params" column, we present the parameter counts of the visual generative model and the significant language encoders for a fair comparison with unified models following [95]. Tables below follow this paradigm.
@@
SD v1.5 [55] 0.9B + 0.1B 2B 0.43 0.97 0.38 0.35 0.76 0.04 0.06
SD v2.1 [55] 0.9B + 0.3B 2B 0.50 0.98 0.51 0.44 0.85 0.07 0.17
SD-XL [48] 2.6B + 0.8B 0.55 0.98 0.74 0.39 0.85 0.15 0.23
SD 3 [15] 8.2B + 2.8B 0.68 0.98 0.84 0.66 0.74 0.40 0.43
DALL-E 2 [50] 4.2B + 1.0B 650M 0.52 0.94 0.66 0.49 0.77 0.10 0.19
FLUX [32] 12B + 2.5B 0.67 0.99 0.85 0.75 0.77 0.22 0.42
@@
LlamaGen [62] 0.8B + 2.9B 60M 0.32 0.71 0.34 0.21 0.58 0.07 0.04
Show-o [81] 1.3B 2B 0.53 0.95 0.52 0.49 0.82 0.11 0.28
Chameleon [65] 34B 1.4B 0.39
Transfusion [95] 7.3B 3.5B 0.63
EMU3 [95] 8B 0.66 0.99 0.81 0.42 0.80 0.49 0.45
LWM [39] 7B 1B 0.47 0.93 0.41 0.46 0.79 0.09 0.15
SEED-X [18] 17B 0.49 0.97 0.58 0.26 0.80 0.19 0.14
Lumos-1 ($352\times 352$) 1.5B 60M 0.601 0.959 0.732 0.375 0.774 0.365 0.400
Lumos-1 ($352\times 352$) 3.6B 60M 0.664 0.953 0.806 0.463 0.806 0.483 0.475
@@ 3.3 Implementation
Architecture. The architecture of Lumos-1 follows Llama [71, 20], meaning that it incorporates RMSNorm [92] and SwiGLU [57] by default. To stabilize training, we integrate query-key normalization (QK-Norm) following Chameleon [65]. We have models of three scales (Lumos-1 0.5B, 1B and 3B), whose architectural details are placed in the Appendix. It is worth noting that we use the 0.5B version for fast ablation studies.
Tokenizers. To unify visual and text token processing, we adopt the discrete version of Cosmos Tokenizer [1] that achieves spatiotemporal compression rates of $8\times 8\times 4$. For text tokens, we retain Chameleon’s text tokenizer [65]. Therefore, Lumos-1’s total codebook size is 129,536, partitioned into 65,536 text tokens and 64,000 visual tokens.
Sequence formatting. The visual tokens and text tokens are interleaved within a sequence [38], with the text tokens specifying metadata, including the text prompt, video resolution, video fps and the number of frames in this video. With this design, we train the model using images and videos of varying aspect ratios without resizing them. The details of the sequence formatting (e.g., organization of special tokens) are placed in the Appendix.
@@
VideoCrafter-I2V [11] 2.6B 20M 82.57 86.31 78.84 91.17 91.31 97.86 98.79 71.68
ConsistI2V [54] 1.3B + 0.3B 10M 84.07 91.91 76.22 95.82 95.95 95.27 98.28 66.92
SEINE [12] 0.9B + 0.1B 10M 84.88 92.39 77.37 96.57 96.80 94.2 97.26 70.97
I2VGen-XL [94] 1.4B + 1.0B 35M 85.28 92.11 78.44 96.48 96.83 94.18 97.09 69.14
CogVideoX [83] 5.6B + 4.8B 86.70 94.79 78.61 97.19 96.74 94.34 96.42 70.01
@@
COSMOS-Video2World [62] 5B + 11B 100M 84.16 92.51 75.81 95.99 97.36 97.12 96.59 59.90
Lumos-1 ($672\times 384\times 25$) 1.5B 10M 84.16 91.80 76.53 96.06 96.58 95.90 96.25 67.06
Lumos-1 ($672\times 384\times 25$) 3.6B 10M 84.72 93.34 76.10 97.42 97.40 97.42 96.91 69.23
Table 3: Performance comparison on VBench-T2V benchmark. We list partial metrics due to space limits.
@@
ModelScopeT2V [73] 1.4B + 0.3B 10M 75.75 78.05 66.54 89.87 95.29 58.57 82.25 81.72 25.67
InstructVideo [91] 1.4B + 0.3B 10M 76.61 81.56 56.81 95.30 96.97 68.01 73.26 77.14 19.91
LaVie [78] 2.5B + 0.5B 25M 77.08 78.78 70.31 91.41 97.47 61.90 91.82 86.39 26.41
OpenSoraPlan V1.3 [35] 2.7B + 13B 70M 77.23 80.14 65.62 97.79 97.24 56.21 85.56 79.30 24.47
LTX-Video [22] 1.9B + 4.8B 80.00 82.30 70.79 96.56 97.20 60.28 83.45 81.45 25.19
CogVideoX [83] 5.6B + 4.8B 81.91 83.05 77.33 96.45 96.71 63.33 85.07 83.03 27.65
@@
CogVideo [26] 9B 5.4M 67.01 72.06 46.83 92.19 96.20 41.03 73.40 79.57 7.70
EMU3 [77] 8B 80.96 84.09 68.43 95.32 97.69 86.17
Lumos-1 ($448\times 256\times 25$) 1.5B 10M 75.78 77.54 68.74 94.97 96.08 54.58 90.08 75.34 24.71
Lumos-1 ($448\times 256\times 25$) 3.6B 10M 78.32 79.52 73.51 95.51 96.50 58.04 90.05 82.00 25.29
[…]
Stage-wise training. Due to the autoregressive nature of Lumos-1, the training of video generation can be categorized into the training of two capabilities: 1) text-to-image and 2) image/images to video. Although the incorporation of AR-DF training substantially ameliorates the imbalanced learning issue, we still observe that the later task is relatively easier. Therefore, a stage-wise training scheme is mandatory to ensure successful video generation training. Concretely, the training commences with dedicated text-to-image training with 256p resolution, and then we perform image-video joint training to train the video generation capability on 256p resolution. Finally, we perform joint training on visual data of 384p resolution.
@@ 4.1 Experimental Details
Training, inference and model evaluation. We train the model from scratch and the basic training hyper-parameters are placed in the Appendix. We adopt the Adam [41] optimizer to optimize the model, set the weight decay to 0.1 and set $\beta_{1}$ and $\beta_{2}$ to $0.9$ and $0.95$. During AR-DF training, the training mask ratio $\rho_{tra}$ is set to $0.7$ following [33]. During AR-DF inference, classier-free guidance (CFG) is utilized by default to enhance the generation quality. The inference mask ratio $\rho_{inf}$ is set to $0.7$ by default. The number of steps to generate one latent frame is set to $N_{steps}=50$ by default. When evaluating on GenEval [19] and VBench [28], the guidance scale is set to 16 if not otherwise specified. Since we train our model only on detailed and descriptive captions, it is mandatory for us to rewrite captions using Qwen 32B [82] when evaluating on GenEval [19]. When evaluating on VBench [28], we use its official long captions by default.
@@ 4.2 Comparison with Other Methods on Visual Generation
Text-to-image generation. We compare Lumos-1 with competitive image generation methods in Tab. 1. Compared with diffusion models, we observe that Lumos-1 outperforms models of similar sizes by a margin (e.g., SD-XL [48]) and is even on par with FLUX [32]. Compared with autoregressive models, we observe that Lumos-1 is on par with EMU3 [77], while being significantly more efficient due to the discrete diffusion inference. Lumos-1 achieves superior results in terms of position and attribute binding, demonstrating excellent language understanding and vision-language alignment even without textual pre-training.
Image-to-video generation. Thanks to the autoregressive nature of Lumos-1, we can perform image-to-video generation by specifying the first frame, although we did not specifically train on this task. Results are listed in Tab. 2. Lumos-1 outperforms the popular VideoCrafter-I2V model and is on par with the leading COSMOS-Video2World model, which uses substantially more data (100M$>$10M) and training resources (10000 H100s$>$48 H20s), demonstrating the promising performance of Lumos-1.
Text-to-video generation. We list the comparison in Tab. 3. Although we utilize discrete tokenizers, Lumos-1 can still be on par with diffusion models like OpenSoraPlan even if we do not rely on a heavy pretrained text encoder. Due to the autoregressive generation nature, the video quality can be ensured by first frame quality, enabling Lumos-1 to excel in object-centric metrics (object class and color).

### v2

@@ 3.1 Spatiotemporal Correlation Injection via MM-RoPE
12:    $\mathcal{C}\leftarrow G_{\phi}(\widetilde{\bm{X}}_{v}^{(t)},AttnMask^{(t)},\mathcal{C})$ $\triangleright$ Store cache for the $t$-th frame
@@
Moreover, for a model that jointly processes text and visual tokens, the interplay between two modalities is significant to ensure the vision-language alignment. However, the range of positions for representing texts or visual data tends to differ. Contemporary visual generation systems are typically trained with extremely long and descriptive captions [5, 70], despite the low latent resolution of visual data (e.g., a video of resolution $448\times 256\times 25$ becomes $56\times 32\times 7$ after a $8\times 8\times 4$ compression). To balance the two modalities, we propose scaling the 3D positions to ensure a balanced learning. Specifically, we empirically scale the latent 3D positions to the RGB space by multiplying the compression ratio, as shown in figure˜4(a). This simple scaling operation, from another perspective, improves the RoPE resolution for visual tokens by slightly accelerating the rotation speed. Experiments in the experiment section demonstrate its efficacy, thereby showing the importance of balancing the two modalities from the RoPE perspective. We acknowledge, however, that given the autoregressive generation nature of videos, this scaling may not be the optimal solution. More advanced and sophisticated solutions are left for future work.
@@ 3.2 Autoregressive Discrete Diffusion Forcing
The most naive generation paradigm (i.e., next-token prediction) suffers from incompatibility with the nature of videos and low generation efficiency, rendering it impractical for autoregressive visual generation. In this work, we resort to discrete diffusion [86, 66] to generate visual content, together with spatially bidirectional and temporally causal token dependency. However, due to the autoregressive nature of Lumos-1, naive random masking (i.e., a globally random mask) or temporally independent masking (i.e., diffusion forcing [9, 63]) both lead to significant loss imbalance, i.e., visual tokens in the later frames tend to have much lower loss. Since the task difficulty of predicting frames with ample history frame context is considerably easier than predicting the first image given a text prompt, the model would lean towards optimizing simpler tasks, leading to a degradation of temporal learning.
Training scheme. To resolve this issue, we build upon the basic nature of videos–spatial information redundancy. The core issue of imbalanced loss during training stems from spatial information leakage. It is worth noting that the original video diffusion transformer [63] that adopts diffusion forcing does not encounter this issue due to its usage of bidirectional dependency. Facing this challenge, we introduce Autoregressive Discrete Diffusion Forcing (AR-DF), which adopts temporal tube masking during the training of AR video generation. Concretely, for every video, we randomly generate a mask pattern for the first frame and then apply this mask pattern repeatedly to later frames in this video. If we denote the multi-modal token sequence $\bm{X}$ composed of text tokens $\bm{X}_{p}$ and visual tokens $\bm{X}_{v}$ and sample a mask ratio $\rho$, the training masking strategy in AR-DF can be formulated as:
@@
$\small\widetilde{\bm{X}}_{v}^{(t)}\;=\;\bm{M}\odot\bm{X}_{v}^{(t)}\;+\;(1-\bm{M})\,\odot\,[\text{MASK}],\quad t=1,\dots,T$ (6)
@@
where $N_{f}$ and $T$ denote the number of tokens in a latent frame and the number of latent frames in $\bm{X}_{v}$; $\bm{X}_{v}^{(t)}$ denotes the visual tokens of the $t$-th frame; $\widetilde{\bm{X}}$ denotes the masked multi-modal token sequence prepared for training; $\odot$ indicates Hadamard multiplication; $\bm{M}$ denotes the mask pattern; $\text{MASK}]$ denotes the mask token. After the preparation of the token sequence, it is fed into the model for processing. To ensure consistency with native LLMs and temporal causality in videos, we adopt a temporal causal mask $AttnMask$ for attention processing, as shown in figure˜4(b). To train the model, we use cross-entropy loss and compute the loss only on masked tokens, denoted as $\mathcal{L}(\widehat{\bm{X}},\bm{X},\bm{M})$ and $\widehat{\bm{X}}$ is the token sequence after model processing. The algorithm is formalized in algorithm˜1.
Inference scheme. After training with AR-DF, the most naive inference scheme (i.e., autoregressively generating video frames) would result in significant frame quality and motion degradation. We observe that this is caused by inconsistent inference with training. During training, later frames consistently have partial observation of history frames, while the inference stage does not align with this observation pattern. Therefore, given a caption, we first generate the first frame by running multiple steps and then randomly replace a pre-defined ratio $\rho_{inf}$ of tokens with the $\text{MASK}]$ token for the generated image. We infer the model with this partially observed image and cache the Keys and Values of this image for swift inference. This process is repeated until the entire video is generated. The algorithm is formalized in algorithm˜2.
Table 2: Performance comparison on GenEval. The “#Params" column shows the parameter counts of the generative model and the language encoder for a fair comparison to unified models following [103]. Gen Rep denotes the generation representation type: continuous or discrete. * denotes results without inference-time scaling. † denotes supervised fine-tuning on small-scale data.
@@
SD v1.5 [59] 0.9B + 0.1B 2B Continuous 0.43 0.97 0.38 0.35 0.76 0.04 0.06
SD-XL [52] 2.6B + 0.8B Continuous 0.55 0.98 0.74 0.39 0.85 0.15 0.23
SD 3 [16] 8.2B + 2.8B Continuous 0.68 0.98 0.84 0.66 0.74 0.40 0.43
DALL-E 2 [54] 4.2B + 1.0B 650M Continuous 0.52 0.94 0.66 0.49 0.77 0.10 0.19
SANA 1.5 [85] 1.6B + 2.6B 50M Continuous 0.66
FLUX [34] 12B + 2.5B Continuous 0.67 0.99 0.85 0.75 0.77 0.22 0.42
Meissonic [3] 1B + 0.4B 210M Discrete 0.54 0.99 0.66 0.42 0.86 0.10 0.22
Muddit [62] 1B + 0.4B 50M Discrete 0.61 0.98 0.72 0.54 0.82 0.19 0.41
UniDisc [66] 1.4B 280M Discrete 0.42 0.92 0.47 0.15 0.67 0.13 0.19
D-DiT [37] 2B Discrete 0.65 0.97 0.80 0.54 0.76 0.32 0.50
MMaDA [90] 8B Discrete 0.63 0.99 0.76 0.61 0.84 0.20 0.37
@@
SEED-X [19] 17B Continuous 0.49 0.97 0.58 0.26 0.80 0.19 0.14
Transfusion [103] 7.3B 3.5B Continuous 0.63
Fluid [17] 10.5B + 4.7B 680M Continuous 0.69 0.96 0.83 0.63 0.80 0.39 0.51
Show-o2 [87] 7B 66M Continuous 0.76 1.00 0.87 0.58 0.92 0.52 0.62
LlamaGen [65] 0.8B + 2.9B 60M Discrete 0.32 0.71 0.34 0.21 0.58 0.07 0.04
Show-o [86] 1.3B 2B Discrete 0.68 0.98 0.80 0.66 0.84 0.31 0.50
Chameleon [69] 34B 1.4B Discrete 0.39
Lumina-mGPT [43] 7B 1.4B Discrete 0.56 0.77 0.27 0.32
Lumina-mGPT 2.0* [88] 7B Discrete 0.73 1.00 0.87 0.49 0.85 0.52 0.62
EMU3 [103] 8B Discrete 0.66 0.99 0.81 0.42 0.80 0.49 0.45
Lumos-1 ($352\times 352$) 1.5B 60M Discrete 0.601 0.959 0.732 0.375 0.774 0.365 0.400
Lumos-1 ($352\times 352$) 3.6B 60M Discrete 0.664 0.953 0.806 0.463 0.806 0.483 0.475
Lumos-1 ($512\times 512$) 1.5B 60M Discrete 0.725 0.984 0.869 0.519 0.862 0.558 0.558
Lumos-1 ($512\times 512$) 3.6B 60M Discrete 0.791 0.991 0.866 0.678 0.840 0.683 0.688
@@ 3.3 Implementation
Architecture. Lumos-1’s architecture follows Llama [75, 21]. To stabilize training, we integrate QK-Norm following Chameleon [69]. We have models of three scales (0.5B, 1B and 3B), whose architectural details are placed in the Appendix.
Tokenizers. We adopt the discrete Cosmos Tokenizer [1] that achieves spatiotemporal compression rates of $8\times 8\times 4$. For text tokens, we retain Chameleon’s text tokenizer. Lumos-1’s total codebook size is 129,536, partitioned into 65,536 text tokens and 64,000 visual tokens.
@@
VideoCrafter-I2V [11] 2.6B 20M Continuous 82.57 86.31 78.84 91.17 91.31 97.86 98.79 71.68
ConsistI2V [58] 1.3B + 0.3B 10M Continuous 84.07 91.91 76.22 95.82 95.95 95.27 98.28 66.92
SEINE [13] 0.9B + 0.1B 10M Continuous 84.88 92.39 77.37 96.57 96.80 94.2 97.26 70.97
I2VGen-XL [102] 1.4B + 1.0B 35M Continuous 85.28 92.11 78.44 96.48 96.83 94.18 97.09 69.14
CogVideoX [91] 5.6B + 4.8B Continuous 86.70 94.79 78.61 97.19 96.74 94.34 96.42 70.01
@@
COSMOS [65] 5B + 11B 100M Continuous 84.16 92.51 75.81 95.99 97.36 97.12 96.59 59.90
VideoMAR [93] 1.4B + 1.5B 0.5M Continuous 84.82 94.02 75.6 97.85 98.38 97.13 97.20 62.34
Lumos-1 ($672\times 384\times 25$) 1.5B 10M Discrete 84.16 91.80 76.53 96.06 96.58 95.90 96.25 67.06
Lumos-1 ($672\times 384\times 25$) 3.6B 10M Discrete 84.72 93.34 76.10 97.42 97.40 97.42 96.91 69.23
Table 4: Performance comparison on VBench-T2V benchmark. We list partial metrics due to space limits. † denotes supervised fine-tuning on small-scale data.
@@
ModelScopeT2V [77] 1.4B + 0.3B 10M Continuous 75.75 78.05 66.54 89.87 95.29 58.57 82.25 81.72 25.67
InstructVideo [99] 1.4B + 0.3B 10M Continuous 76.61 81.56 56.81 95.30 96.97 68.01 73.26 77.14 19.91
LaVie [82] 2.5B + 0.5B 25M Continuous 77.08 78.78 70.31 91.41 97.47 61.90 91.82 86.39 26.41
OpenSoraPlan V1.3 [40] 2.7B + 13B 70M Continuous 77.23 80.14 65.62 97.79 97.24 56.21 85.56 79.30 24.47
LTX-Video [23] 1.9B + 4.8B Continuous 80.00 82.30 70.79 96.56 97.20 60.28 83.45 81.45 25.19
CogVideoX [91] 5.6B + 4.8B Continuous 81.91 83.05 77.33 96.45 96.71 63.33 85.07 83.03 27.65
@@
CogVideo [27] 9B 5.4M Continuous 67.01 72.06 46.83 92.19 96.20 41.03 73.40 79.57 7.70
NOVA [15] 0.6B + 2.8B 20M Continuous 80.12 80.39 79.05 92.00
EMU3 [81] 8B Discrete 80.96 84.09 68.43 95.32 97.69 86.17
Lumos-1 ($672\times 384\times 25$) 1.5B 10M Discrete 76.34 78.27 68.65 96.04 96.29 56.16 90.47 77.49 25.07
Lumos-1 ($672\times 384\times 25$) 3.6B 10M Discrete 78.32 79.52 73.51 95.51 96.50 58.04 90.05 82.00 25.29
Lumos-1 ($672\times 384\times 25$) 1.5B 10M Discrete 78.17 79.60 72.45 95.36 96.08 57.21 93.39 77.13 25.55
Lumos-1 ($672\times 384\times 25$) 3.6B 10M Discrete 78.90 79.83 75.21 96.78 96.60 61.92 94.38 81.14 25.57
Stage-wise training. Due to the autoregressive nature of Lumos-1, the training of video generation can be categorized into training two capabilities: 1) text-to-image and 2) image/images to video. Although the incorporation of AR-DF training substantially ameliorates the imbalanced learning issue, we still observe that the later task is relatively easier. Therefore, a stage-wise training scheme is mandatory to ensure successful video generation training. Our training consists of three pre-training stages and a supervised fine-tuning (SFT) stage. The pre-training stages focus on learning the generation capabilities from a vast amount of data: 1) text-to-image training at 256p, 2) joint image-video training at 256p, and 3) joint fine-tuning at 384p. The final SFT stage focuses on small-scale fine-tuning on data of high quality. We use around 100k images and 20k videos. This process teaches the model to more precisely follow instructions and adhere to specific aesthetic styles. We find that both models trained on 256p or 384p data can be boosted by simple SFT.
Details on sequence formatting and GPU memory friendly implementation are placed in the Appendix.
@@ 4.1 Experimental Details
Training, inference and model evaluation. We train the model from scratch and the basic training hyper-parameters are placed in the Appendix. During AR-DF training, the training mask ratio $\rho_{tra}$ is set to $0.7$ following [36]. During AR-DF inference, classier-free guidance (CFG) is utilized by default to enhance the generation quality. The inference mask ratio $\rho_{inf}$ is set to $0.7$ by default. The number of steps to generate one latent frame is set to $N_{steps}=50$ by default. When evaluating on GenEval [20] and VBench [30], the guidance scale is set to 16 if not otherwise specified. Since we train our model only on detailed and descriptive captions, it is mandatory for us to rewrite captions using Qwen 32B [89] when evaluating on GenEval [20]. When evaluating on VBench [30], we use its official long captions by default.
@@ 4.2 Comparison with Other Methods on Visual Generation
Text-to-image generation. We compare Lumos-1 with competitive image generation methods in Tab.˜2. Compared with diffusion models, we observe that Lumos-1 outperforms competitive models (e.g., SD-XL [52] and FLUX [34]) using continuous or discrete tokenizers by a clear margin, even if Lumos-1 only uses a discrete tokenizer. Compared with autoregressive models, we observe that Lumos-1 outperforms EMU3 [81], Fluid [17] and Lumina-mGPT series [88, 43], while using substantially fewer training images and retaining a small model. Lumos-1 achieves superior results in terms of position and attribute binding, demonstrating excellent language understanding and vision-language alignment even without textual pre-training. We also observe that with a small-scale supervised fine-tuning using high-quality data, the performance can be largely boosted on GenEval, especially on challenging metrics like position and attribute binding.
Image-to-video generation. Thanks to the autoregressive nature of Lumos-1, we can perform image-to-video generation by specifying the first frame, although we did not specifically train on this task. Results are listed in Tab.˜3. Lumos-1 outperforms the popular VideoCrafter-I2V model and is on par with the leading COSMOS-Video2World model, which uses substantially more data (100M>>10M) and training resources (10000 H100s>>48 H20s), demonstrating the promising performance of Lumos-1.
Table 5: Comparison with different training methods for AR video generation. Results are measured on GenEval, Vbench-Overall Consistency (OC), and Vbench-Imaging Quality (IQ).
Training Methods GenEval Vbench-OC Vbench-IQ
Random masks 0.593 0.232 0.424
Linear decay loss 0.588 0.228 0.410
Step decay loss 0.584 0.234 0.464
Diffusion forcing masks 0.590 0.241 0.540
AR-DF 0.591 0.249 0.559
Text-to-video generation. We list the comparison in Tab.˜4. Although Lumos-1 utilizes discrete tokenizers and does not rely on a heavy pretrained text encoder, it can still be on par with diffusion models like OpenSoraPlan. Due to the autoregressive nature, the video quality can be ensured by first frame quality, enabling Lumos-1 to excel in object-centric metrics (object class and color).